# Stability Analysis of Cosmological Perturbations in the Bumblebee Model: Parameter Constraints and Gravitational Waves

### v2

@@ VI.2 Gauge I: $\alpha=0,\;E=0$
The action (102) is an effective action that contains only the dynamical variables $\psi_{3}$ and $\phi_{m}$. The free of ghost instability requires
$\frac{3(1-\xi b_{t}^{2})}{\kappa b_{t}^{2}}a^{3}H^{2}$ $>$ $0,$ (103)
@@
Since $|\xi b_{t}^{2}|\ll 1$, $a>0$, and $\kappa>0$, the perturbation $\psi_{3}$ is free of ghost instability. Equations (18) and (5), together with the conditions $\overline{\rho}_{m}>0$ and $\overline{p}_{m}\geq 0$, imply that $\overline{J}>0$ and $\overline{\rho}_{m,n}>0$. Thus, the perturbation $\phi_{m}$ is also ghost-free.
@@
$\frac{3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)}{3\kappa(1-\xi b_{t}^{2})}$ $>$ $0,$ (105)
@@
The first condition (105) is obviously satisfied, as will be shown in Eq. (109) below. Furthermore, the condition (106) implies $\overline{p}_{m,n}>0$. This means that the greater the particle number density, the higher the pressure $\overline{p}_{m}$ contributed by the matter, which is a result consistent with physical intuition.
@@
$c_{s}^{2}=\frac{3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)}{3\kappa(1-\xi b_{t}^{2})}=1+\frac{\kappa-2\xi}{3\kappa}(-\xi b_{t}^{2})+\mathcal{O}\big((\xi b_{t}^{2})^{2}\big).$ (109)
@@
Avoiding tachyonic instability in the vector perturbations requires $\xi\leq 0$, as shown in Eq. (77). Consequently, the propagation speed of the scalar perturbations exceeds 1, which results from the violation of Lorentz symmetry.
VI.3 Gauge II: $\alpha=0,\varphi_{h}=0$
Similar to the previous analysis, we analyze the stabilities of the scalar perturbations under gauge condition $\alpha=0,\varphi_{h}=0$ in Fourier space. After eliminating all variables without kinetic terms, the effective action can be written as
$S^{(2)}_{s}=\int dtd^{3}x$ $\Bigg[$ $\frac{(1-\xi b_{t}^{2})aH^{2}}{2(1-\xi b_{t}^{2})H^{2}+\kappa b_{t}^{2}\dot{H}}(k\dot{\psi}_{7})^{2}+\frac{3\xi^{2}b_{t}^{4}a^{3}}{4\kappa(1-\xi b_{t}^{2})}(\dot{\psi}_{5})^{2}+\frac{a^{2}}{2\overline{J}}\overline{\rho}_{m,n}(\dot{\phi}_{m})^{2}+f(t)b_{t}^{2}ak^{2}\psi_{5}^{2}$ (110)
$-\frac{\overline{\rho}_{m,nn}}{2a^{3}}k^{2}\left(\frac{\phi_{m}}{k}\right)^{2}+\frac{b_{t}\big(\xi b_{t}^{2}(2\xi-3\kappa)+2\kappa\big)}{\kappa\big(2(1-\xi b_{t}^{2})H^{2}+\kappa b_{t}^{2}\dot{H}\big)}aH^{2}k\psi_{5}(k\dot{\psi}_{7})+M_{2}\left(k\psi_{7},\frac{1}{k}\phi_{m},\psi_{5}\right)\;\;\Bigg],$
where $M_{2}$ includes all terms which are of order smaller than $k^{2}$. The specific form of $f(t)$ is
$f(t)=\frac{(1-\xi b_{t}^{2})\big(\kappa(2-3\xi b_{t}^{2})^{2}+2\xi^{2}b_{t}^{2}(1-2\xi b_{t}^{2})\big)H^{2}-\xi^{2}b_{t}^{4}\big(3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)\big)\dot{H}}{4\kappa(1-\xi b_{t}^{2})^{2}\big(2(1-\xi b_{t}^{2})H^{2}+\kappa b_{t}^{2}\dot{H}\big)}.$ (111)
During the derivation of this action, we successively introduce new variables
$\psi_{4}\equiv E+\frac{\xi b_{t}^{2}}{1-\xi b_{t}^{2}}\phi_{h}-\frac{\xi b_{t}}{1-\xi b_{t}^{2}}\phi_{b},\qquad\psi_{5}\equiv\phi_{h}-\frac{1}{b_{t}}\phi_{b},\qquad\psi_{6}\equiv\varphi_{b}-\frac{b_{t}}{2H}\psi_{4},$ (112)
$\psi_{7}\equiv\psi_{6}+\frac{\kappa b_{t}a\overline{\rho}_{m,n}}{18(1-\xi b_{t}^{2})a^{6}H^{2}(3H\dot{H}+\ddot{H})-27\kappa\overline{J}^{2}H^{3}\overline{\rho}_{m,nn}+2(1-\xi b_{t}^{2})a^{2}H(k^{2}-3a^{2}\dot{H})k^{2}}k^{2}\phi_{m}.$
In the small-scale limit, varying $S^{(2)}_{s}$ with respect to $\psi_{7}$, we can obtain the constraint (analogous to the treatment of the vector perturbations)
$k\dot{\psi}_{7}\approx-\frac{b_{t}\big(2\kappa+\xi b_{t}^{2}(2\xi-3\kappa)\big)k}{2\kappa(1-\xi b_{t}^{2})}\psi_{5}.$ (113)
Using this constraint, an effective action containing only the dynamical variables in small-scale limit is obtained
$S^{(2)}_{s}\approx\int dtd^{3}x\left[\frac{3\xi^{2}b_{t}^{4}}{4\kappa(1-\xi b_{t}^{2})}a^{3}(\dot{\psi}_{5})^{2}+\frac{a^{2}}{2\overline{J}}\overline{\rho}_{m,n}\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}-\frac{\xi^{2}b_{t}^{4}\left(2\xi b_{t}^{2}(\xi-2\kappa)+3\kappa\right)}{4\kappa^{2}(1-\xi b_{t}^{2})^{2}}ak^{2}\psi_{5}^{2}-\frac{1}{2a^{3}}\overline{\rho}_{m,nn}k^{2}\left(\frac{\phi_{m}}{k}\right)^{2}\right].$ (114)
Varying this action with respect to $\psi_{5}$ and $\phi_{m}$, respectively, we obtain their dispersion relations in the small-scale limit
$w_{\psi_{5}}^{2}-\frac{3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)}{3\kappa(1-\xi b_{t}^{2})}\overline{g}^{ij}k_{i}k_{j}$ $=$ $0,$ (115)
Clearly, the dispersion relations (115) and (116) are identical to those in the case with the gauge condition $\alpha=0,E=0$. Moreover, the conditions for Laplacian stability and ghost stability yield identical results under both gauge choices.
VI.4 Gauge III: $\alpha=0,\varphi_{b}=0$
In this subsection, we focus on the gauge condition $\alpha=0,\varphi_{b}=0$. In Fourier space, after eliminating all variables without kinetic terms, the effective action can be written as
$S^{(2)}_{s}=\int dtd^{3}x$ $\Bigg[$ $\frac{b_{t}^{2}(1-\xi b_{t}^{2})}{8(1-\xi b_{t}^{2})H^{2}+4\kappa b_{t}^{2}\dot{H}}a(k\dot{\psi}_{8})^{2}+\frac{3\xi^{2}b_{t}^{2}}{4\kappa(1-\xi b_{t}^{2})}a^{3}(\dot{\psi}_{9})^{2}+\frac{a^{2}}{2\overline{J}}\overline{\rho}_{m,n}\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}+f(t)ak^{2}\psi_{9}$ (117)
$-\frac{\overline{\rho}_{m,nn}}{2a^{3}}k^{2}\left(\frac{\phi_{m}}{k}\right)^{2}+\frac{b_{t}\big(2\kappa+\xi b_{t}^{2}(2\xi-3\kappa)\big)}{2\kappa\big(2(1-\xi b_{t}^{2})H^{2}+\kappa b_{t}^{2}\dot{H}\big)}aHk\psi_{9}(k\dot{\psi}_{8})+M_{3}\left(k\psi_{8},\frac{1}{k}\phi_{m},\psi_{9}\right)\;\;\Bigg].$
Here, $M_{3}$ includes all terms which are of order smaller than $k^{2}$. During the derivation of this action, we successively introduce the new variables
$\psi_{8}\equiv E+\frac{\xi b_{t}^{2}}{1-\xi b_{t}^{2}}\phi_{h}-\frac{\xi b_{t}}{1-\xi b_{t}^{2}}\phi_{b}-\frac{\kappa\overline{\rho}_{m,n}}{(1-\xi b_{t}^{2})(k^{2}-3a^{2}H^{\prime})a}\phi_{m},\qquad\psi_{9}\equiv\phi_{b}-b_{t}\phi_{h}.$ (118)
In the small-scale limit, varying $S^{(2)}_{s}$ with respect to $\psi_{8}$, we can obtain the constraint (analogous to the treatment of the vector perturbations)
$k\dot{\psi}_{8}\approx-\frac{\big(2\kappa+\xi b_{t}^{2}(2\xi-3\kappa)\big)k}{\kappa b_{t}(1-\xi b_{t}^{2})}H\psi_{9}.$ (119)
with this constraint, an effective action which contains only dynamical variables in small-scale limit is obtained
$S^{(2)}_{s}\approx\int dtd^{3}x\left[\frac{3\xi^{2}b_{t}^{4}}{4\kappa(1-\xi b_{t}^{2})}a^{3}(\dot{\psi}_{9})^{2}+\frac{a^{2}}{2\overline{J}}\overline{\rho}_{m,n}\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}-\frac{\xi^{2}b_{t}^{2}\big(2\xi b_{t}^{2}(\xi-2\kappa)+3\kappa\big)}{4\kappa^{2}(1-\xi b_{t}^{2})^{2}}ak^{2}\psi_{9}^{2}-\frac{1}{2a^{3}}\overline{\rho}_{m,nn}k^{2}\left(\frac{\phi_{m}}{k}\right)^{2}\right].$ (120)
It is evident that the dispersion relations are identical to those found for both the gauge conditions $\alpha=0,\varphi_{h}=0$ and $\alpha=0,E=0$. Furthermore, the stability conditions are the same as those given in the case with the gauge condition $\alpha=0,E=0$ and in the case with the gauge condition $\alpha=0,\varphi_{h}=0$.
Until now, we have analyzed the dispersion relations and the system’s stability under three different gauge choices. We found that they yield identical results, which is expected since gauge degrees of freedom do not contribute to physical effects. The matter sound speed is given by $\sqrt{\overline{p}_{m,n}/\overline{\rho}_{m,n}}$. The propagation speed of the scalar perturbations is greater than 1, due to the presence of the Lorentz-violating parameter $\xi b_{t}^{2}$. Since $|\xi b_{t}^{2}|\ll 1$, the scalar perturbations in the Bumblebee model are free of ghost instability and Laplacian instability.

### v3

@@ VI.2 Effective action and stability conditions
where $F_{1}(t,k)=H^{2}\big(1-(\xi+\sigma)b_{t}^{2}\big)\big(k^{2}-3a^{2}\dot{H}\big)$ and all non-kinetic terms are contained in $\cdots$”. The action (99) is an effective action containing only the dynamical variables $\psi_{3}$, $\phi_{m}$, and $\psi_{2}$, while the variables $\phi_{h}$, $\varphi_{h}$, $\phi_{b}$, $\varphi_{b}$, $\varphi_{m}$, and $\phi_{\ell}$ depend on them. From the structure of the kinetic terms in the action, one finds that the free of ghost instability requires
$\frac{3(\xi+2\sigma)^{2}b_{t}^{2}a^{3}}{4\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}>0,$ (100)
$\frac{a^{2}\overline{\rho}_{m,n}k^{2}}{2\overline{J}(k^{2}-3a^{2}\dot{H})}>0,$ (101)
$\frac{aF_{1}(t,k)}{\kappa b_{t}^{2}\dot{H}k^{2}+2F_{1}(t,k)}>0.$ (102)
Using the assumes $\overline{\rho}_{m}>0$ and $\overline{p}_{m}>0$, the condition $\overline{\rho}_{m,n}>0$ can be obtained by Eq. (5). Since $a>0$, $\kappa>0$, $\overline{J}>0$, and $\dot{H}<0$, the condition (101) is satisfied, while the conditions (100) and (102) reduce to
$1-(\xi+\sigma)b_{t}^{2}>0,$ (103)
$\frac{1}{k^{2}}>\frac{1}{a^{2}}\Big(\frac{\kappa b_{t}^{2}}{6\big(1-(\xi+\sigma)b_{t}^{2}\big)H^{2}}+\frac{1}{3\dot{H}}\Big).$ (104)
It is not difficult to see that the condition (103) holds true, since $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$. According to the condition (104), the absence of ghost instability in the scalar perturbations may depend on the wavenumber $k$, and thus be frequency-dependent. If $\tfrac{\kappa b_{t}^{2}}{2\big(1-(\xi+\sigma)b_{t}^{2}\big)H^{2}}+\tfrac{1}{\dot{H}}>0$, the condition (104) fails for large $k$ (i.e., in the high-frequency/short-wavelength perturbation), leading to a ghost instability. A physically viable theory of gravity must avoid ghost instability at all perturbation frequencies. Therefore, to ensure ghost-free perturbations for all $k$, we must enforce the stronger requirement that $\tfrac{\kappa b_{t}^{2}}{2\big(1-(\xi+\sigma)b_{t}^{2}\big)H^{2}}+\tfrac{1}{\dot{H}}\leq 0$. This inequality can be simplified to
$\frac{1-(\xi+\sigma)b_{t}^{2}}{b_{t}^{2}}\geq-\frac{\kappa\dot{H}}{2H^{2}}.$ (105)
Once astronomical observations determine the values of $H^{2}$ and $\dot{H}$, substituting them into this equation will impose a constraint on the parameters $\xi$, $\sigma$, and $b_{t}$. Further applying the condition $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$ yields an upper bound for $b_{t}^{2}$: $b_{t}^{2}\lesssim-\tfrac{2H^{2}}{\kappa\dot{H}}$.
Following the same procedure, we analyze the stability of scalar perturbations in Fourier space under two distinct gauge conditions: (i) $\alpha=0$, $\varphi_{h}=0$, and (ii) $\alpha=0$, $\varphi_{b}=0$.
Gauge condition $\alpha=0,\;\varphi_{h}=0$
After imposing the gauge condition $\alpha=0$, $\varphi_{h}=0$ and eliminating all variables without kinetic terms in the action (83), we obtain an effective action in Fourier space
$S^{(2)}_{s}=\int dtd^{3}x\left[\frac{3(\xi+2\sigma)^{2}b_{t}^{4}a^{3}}{4\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}\dot{\psi}_{5}^{2}+\frac{a^{2}\overline{\rho}_{m,n}k^{2}}{2\overline{J}(k^{2}-3a^{2}\dot{H})}\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}+\frac{aF_{1}(t,k)}{\kappa b_{t}^{2}\dot{H}k^{2}+2F_{1}(t,k)}(k\dot{\psi}_{7})^{2}+\cdots\right].$ (106)
Here, all non-kinetic terms are contained in $\cdots$”. During the derivation of this action, we successively introduce new variables
$\psi_{4}\equiv E+\frac{(\xi+2\sigma)b_{t}^{2}}{1-(\xi+\sigma)b_{t}^{2}}\phi_{h}-\frac{(\xi+2\sigma)b_{t}}{1-(\xi+\sigma)b_{t}^{2}}\phi_{b},\qquad\psi_{5}\equiv\phi_{h}-\frac{1}{b_{t}}\phi_{b},$ (107)
$\psi_{6}\equiv\varphi_{b}-\frac{b_{t}}{2H}\psi_{4},\qquad\psi_{7}\equiv\psi_{6}+\frac{\kappa b_{t}\overline{\rho}_{m,n}}{2aH\big(1-(\xi+\sigma)b_{t}^{2}\big)(k^{2}-3a^{2}\dot{H})}k^{2}\phi_{m}.$
Comparing the actions in Eqs. (99) and action (106) shows that the scalar perturbations in the gauge $\alpha=0$, $E=0$ satisfy the same ghost-free conditions as those in the gauge $\alpha=0$, $\varphi_{h}=0$.
Gauge condition $\alpha=0,\;\varphi_{b}=0$
Under the gauge condition $\alpha=0$, $\varphi_{b}=0$, we eliminate the non-dynamical variables from the action in Eq. (83) to obtain the following effective action in Fourier space
$S^{(2)}_{s}=\int dtd^{3}x\left[\frac{3(\xi+2\sigma)^{2}b_{t}^{2}a^{3}}{4\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}\dot{\psi}_{9}^{2}+\frac{a^{2}\overline{\rho}_{m,n}k^{2}}{2\overline{J}(k^{2}-3a^{2}\dot{H})}\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}+\frac{b_{t}^{2}}{4H^{2}}\frac{aF_{1}(t,k)}{\kappa b_{t}^{2}\dot{H}k^{2}+2F_{1}(t,k)}(k\dot{\psi}_{8})^{2}+\cdots\right].$ (108)
Here, all non-kinetic terms are contained in $\cdots$”. In deriving this action, we successively introduce a series of new variables
$\psi_{8}\equiv E+\frac{(\xi+2\sigma)b_{t}^{2}}{1-(\xi+\sigma)b_{t}^{2}}\phi_{h}-\frac{(\xi+2\sigma)b_{t}^{2}}{1-(\xi+\sigma)b_{t}^{2}}\phi_{b}-\frac{\kappa\overline{\rho}_{m,n}}{\big(1-(\xi+\sigma)b_{t}^{2}\big)\big(k^{2}-3a^{2}\dot{H}\big)a}\phi_{m},\qquad\psi_{9}\equiv\phi_{b}-b_{t}\phi_{h}.$ (109)
A comparison of this action (108) with the actions in Eqs. (99) and (106) reveals that its ghost-free conditions are identical in the gauge $\alpha=0$, $E=0$ and the gauge $\alpha=0$, $\varphi_{h}=0$.
We have analyzed the stability of scalar perturbations under three different gauge choices and found identical results, as expected, given that gauge degrees of freedom do not affect physical observables. The effective actions in Eqs. (99), (106), and (108) contain only dynamical variables, which confirms that the Bumblebee model contains two scalar modes, in addition to the one in the matter sector. Regarding ghost instability, the condition (Eq. (104)) for its absence can, in principle, depend on frequency (or wavenumber). However, under the small-parameter assumption $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$, the condition (105), which reduces to $b_{t}^{2}\lesssim\tfrac{2H^{2}}{-\kappa\dot{H}}$, ensures that the scalar perturbations are ghost-free at all frequencies.
Note that an examination of the key Eqs. (95), (96), (98), (99), (106), and (108) suggests that the aforementioned stability conditions may no longer hold under the parameter condition $\sigma=-\frac{1}{2}\xi$. However, under the condition $\sigma=-\frac{1}{2}\xi$, the background equation (25) reduces to $H^{2}=-2\kappa\overline{V}_{,b^{2}}/(3\xi)=constant$, which is inconsistent with astronomical observations. Consequently, this case is excluded from our analysis. It should be noted that a time-dependent background value $b_{t}$ does not generally yield a constant Hubble parameter under the condition $\sigma=-\frac{1}{2}\xi$. This point is discussed in Ref. van de Bruck et al. (2025), where it was shown that applying this condition results in only one scalar mode, in addition to the one present in the matter sector.
VI.3 The small-scale limit
In the previous subsection, we studied the ghost-free conditions for the scalar perturbations in the Bumblebee model. We now turn to the stability and propagation characteristics of GWs in the small-scale limit ($k\rightarrow\infty$). Since the physics is gauge-invariant, we perform this analysis specifically in the gauge condition $\alpha=0,E=0$.
In the small-scale limit, retaining only the $k^{2}$-order terms, the action (99) simplifies to
$S^{(2)}_{s}\approx\int dtd^{3}x$ $\Bigg[$ $\frac{3(\xi+2\sigma)^{2}b_{t}^{2}a^{3}}{4\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}(\dot{\psi}_{3})^{2}+\frac{a^{2}\overline{\rho}_{m,n}}{2\overline{J}}\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}-\frac{F_{2}(t)}{4\kappa}ak^{2}\psi_{3}^{2}-\frac{\overline{\rho}_{m,nn}}{2a^{3}}k^{2}\left(\frac{\phi_{m}}{k}\right)^{2}$ (110)
$+\frac{\big(1-(\xi+\sigma)b_{t}^{2}\big)aH^{2}}{2\big(1-(\xi+\sigma)b_{t}^{2}\big)H^{2}+\kappa b_{t}^{2}\dot{H}}(k\dot{\psi}_{2})^{2}-\frac{\kappa(2-3\xi b_{t}^{2})+2(\xi+2\sigma)\xi b_{t}^{2}}{\kappa\big(\kappa\dot{H}b_{t}^{2}+2\big(1-(\xi+\sigma)b_{t}^{2}\big)H^{2}\big)}aH^{2}k\psi_{3}(k\dot{\psi}_{2})\;\Bigg],$
where the specific form of $F_{2}(t)$ is
$F_{2}(t)$ $=$ $\Big(-\big(1-(\xi+\sigma)b_{t}^{2}\big)\big(\kappa(2-3\xi b_{t}^{2})^{2}+2b_{t}^{2}(\xi-6\sigma)(\xi+2\sigma)\big)\big(1-(2\xi+\sigma)b_{t}^{2}\big)H^{2}$ (111)
$+b_{t}^{4}\dot{H}(\xi+2\sigma)^{2}\big(2\xi^{2}b_{t}^{2}+\kappa\big(3-(4\xi+3\sigma)b_{t}^{2}\big)\big)\Big)/\Big(\big(1-(\xi+\sigma)b_{t}^{2}\big)^{2}\big(\kappa b_{t}^{2}\dot{H}+2\big(1-(\xi+\sigma)b_{t}^{2}\big)H^{2}\big)\Big).$
In this action, it is $\dot{\psi}_{2}$, and not $\psi_{2}$ itself, that contributes to the $k^{2}$-order terms associated with the variable $\psi_{2}$. Varying the action (110) with respect to $\psi_{2}$ yields the following equation of motion (analogous to the treatment of the vector perturbations)
$k\dot{\psi}_{2}\approx\frac{\kappa(2-3\xi b_{t}^{2})+2\xi b_{t}^{2}(\xi+2\sigma)}{2\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}k\psi_{3}.$ (112)
Here, we consider only the dynamical effects and have accordingly simplified the expression. Substituting this constraint into the action (110), an effective action that contains only the dynamical variables is obtained
$S^{(2)}_{s}\approx\int dtd^{3}x\left[\frac{3(\xi+2\sigma)^{2}b_{t}^{2}a^{3}}{4\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}\left((\dot{\psi}_{3})^{2}-c_{s}^{2}\frac{k^{2}}{a^{2}}\psi_{3}^{2}\right)+\frac{a^{2}\overline{\rho}_{m,n}}{2\overline{J}}\left(\left(\frac{\dot{\phi}_{m}}{k}\right)^{2}-\frac{\overline{p}_{m,n}}{\overline{\rho}_{m,n}}\frac{k^{2}}{a^{2}}\left(\frac{\phi_{m}}{k}\right)^{2}\right)\right],$ (113)
where $c_{s}^{2}\equiv\big(3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)-3\kappa\sigma b_{t}^{2}\big)/\big(3\kappa(1-(\xi+\sigma)b_{t}^{2})\big)$. It is straightforward to see that the variables $\phi_{h}$, $\varphi_{h}$, $\phi_{b}$, $\varphi_{b}$, $\varphi_{m}$, and $\phi_{\ell}$ depend on $\psi_{2}$, $\psi_{3}$, and $\phi_{m}$, as shown in Eqs. (92)-(98). The action (110) indicates that $\psi_{2}$ is a cyclic coordinate in the small-scale limit, and the constraint (112) implies that $\dot{\psi}_{2}$ depends on $\psi_{3}$.
The action (113) is an effective action that contains only the dynamical variables $\psi_{3}$ and $\phi_{m}$. The free of ghost instability in the small-scale limit requires
$\frac{3(\xi+2\sigma)^{2}b_{t}^{2}a^{3}}{4\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}$ $>$ $0,$ (114)
@@ VI.3 The small-scale limit
Since $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$, $a>0$, and $\kappa>0$, the perturbation $\psi_{3}$ is free of ghost instability in the small-scale limit. For the variable $\phi_{m}$, since $\overline{J}>0$ and $\overline{\rho}_{m,n}>0$, it is also ghost-free in the small-scale limit.
@@
$\frac{3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)-3\kappa\sigma b_{t}^{2}}{3\kappa(1-(\xi+\sigma)b_{t}^{2})}$ $>$ $0,$ (116)
@@
The first condition (116) is obviously satisfied since $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$, as will be shown in Eq. (120) below. Furthermore, the condition (117) implies $\overline{p}_{m,n}>0$. This means that the greater the particle number density, the higher the pressure $\overline{p}_{m}$ contributed by the matter, which is a result consistent with physical intuition.
@@
$c_{s}^{2}=\frac{3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)-3\kappa\sigma b_{t}^{2}}{3\kappa(1-(\xi+\sigma)b_{t}^{2})}=1+\frac{\kappa-2\xi}{3\kappa}(-\xi b_{t}^{2})+\mathcal{O}\big((\xi b_{t}^{2})^{n}(\sigma b_{t}^{2})^{m}\big),$ (120)
@@
where $n+m\geq 2$. Avoiding tachyonic instability in the vector perturbations requires $\xi\leq 0$, as shown in Eq. (76). Consequently, the propagation speed of the scalar perturbations exceeds 1, which results from the violation of Lorentz symmetry. If the Lorentz-violating parameter $\xi b_{t}^{2}$ equals zero, the speed will return to 1.
In conclusion, in the small-scale limit, the matter sound speed is given by $\sqrt{\overline{p}_{m,n}/\overline{\rho}_{m,n}}$. The propagation speed of scalar GWs is greater than 1, a direct consequence of Lorentz violation through the parameter $\xi b_{t}^{2}$. Indeed, this superluminality vanishes when $\xi b_{t}^{2}=0$, in which case the scalar GW speed equals 1 for any $\sigma$. Finally, the scalar modes in the Bumblebee model are free of both ghost and Laplacian instabilities, provided that $b_{t}^{2}\lesssim\tfrac{2H^{2}}{-\kappa\dot{H}}$, $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$, and $\overline{p}_{m,n}>0$.